the perimeter of a standard-size rectangular rug is 44 ft. the length is 2 ft longer than the width. find the dimensions
Answer by mananth(16946)

let width be w
length = w+2
..
Perimeter = 2*(l+w)
2(w+w+2)= 44
2(2w+2)=44
4w+4=44
4w=44-4
4w=40
/4
w=10 the width
length = 10+2 = 12
..
12+12+10+10=44
....
m.ananth@hotmail.ca